# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

### v2

@@ 4.1 Framework
To satisfy the desiderata discussed above, we propose a versatile framework for autoregressive diffusion called ACDiT. For generality, ACDiT runs block-wise autoregression instead of token-wise autoregression. We identify there are two kinds of blocks: the clean blocks $c_{i}$ and the noise blocks $n_{i}$, where $n_{i}$ is corrupted from $c_{i}$. ACDiT learns several conditional distribution $p(c_{i}|c_{<i})$ factorized with Eq. 1, where $p(c_{i}|c_{<i})$ is optimized by learning the conditional noise prediction network $\epsilon_{\theta}(n_{i}^{(t)};t,c_{<i})$ with Eq. 2. The final training objective is:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\sim U[0,1],\epsilon\sim\mathcal{N}(0,I)}\sum_{i=1}^{n}||\epsilon_{\theta}(n_{i}^{(t)};t,c_{<i})-\epsilon^{(t)}||_{2}.$ (3)
Given the aforementioned desiderata, we can conceptualize all $n_{i}$ and all $c_{i}$ as occupying separate positions, effectively transforming the dependency structure into an attention pattern between different positions. We designate this attention pattern as the Skip-Causal Attention Mask (SCAM), which is visually represented in Figure 2(a). The figure elucidates that $n_{i}$ attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself, while $c_{i}$ also attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself. In training, for both simplicity and efficiency, we can group the attention mechanisms as illustrated in the right matrix of Figure 2(a). Suppose the number of blocks is $N$, then the unmasked positions form two triangular matrices of side length $N-1$, complemented by a diagonal matrix of side length $N$.
During the inference phase, each autoregressive step executes a conditional diffusion process for $n_{i}$ based on $\{c_{j}|j=0,...,i-1\}$’s KV-Cache. Upon finishing denoising, it is appended to the clean sequence as $c_{i}$ followed by the maximal noise version of the next block $n_{i+1}$. The key-value tensor will be computed for these two blocks, and the key-value tensor of the clean block $c_{i}$ will be kept in KV-Cache. All noise-corrupted version of $n_{i}$ is disregarded. The process is visualized in Figure 2(b). A three-dimensional view of our method is presented in Figure 2(c). By interpolating between full-sequence diffusion and autoregressive paradigms, ACDiT enjoys flexibility and expressivity, enabling it to generate video of any length utilizing the latest long-context techniques developed for language models.
@@ 4.2 Positional Encoding
ACDiT is designed to be versatile, capable of handling one-, two-, three-, or even higher-dimensional data, including but not limited to text (1D), images (2D), and video (3D). For any given dimension of data, the position of that data is a critical attribute that must be made known to the model. This positional awareness enables the model to contextualize its current focus relative to historical data. In the domain of textual data, the Rotary Position Embedding (RoPE) [66] has gained widespread adoption as an effective relative positional encoding method. To address the challenges posed by multi-dimensional positional indices, we introduce RoPE-ND, a natural extension of RoPE.
For a token of a $D$ dimensional data, its positional index is $[m_{1},m_{2},...,m_{D}]$. Given query and key vectors in Transformer’s attention module, we partition the hidden dimension into $D$ segments. It is imperative that each segment’s hidden dimension be an even number. For each segment $j$, we apply a RoPE with a specific base $b_{j}$, as defined in the following equation:
@@
In this formulation, each $\mathbf{R}^{d_{j}}_{\Theta_{j},m_{j}}$ represents a $d_{j}$-dimensional rotary matrix [^2] See a detailed explanation in the Equation 15 in  [66]. with rotation angles $\Theta_{j}=\{\theta_{ji}=b_{j}^{-2(i-1)/d_{j}},i\in[1,2,\cdots,\frac{d_{j}}{2}]\}$. The base $b_{j}$ is empirically determined as $100\lceil\frac{8L_{j}}{100\pi}\rceil$, where $L_{j}$ denotes the maximum position index in data dimension $j$. This formulation ensures that the highest wavelength of RoPE is approximately eight times the maximum position, thereby mitigating rapid decay in long-term dependencies.
It is worth noting that ACDiT inherently supports length extrapolation [66], although a comprehensive exploration of this capability falls beyond the scope of the present work.
4.3 Efficiency Analysis and Block Size Choice
In this section, we provide a brief analysis on the computational efficiency of ACDiT in terms of floating-point operations (FLOPS). We first assume that denoising each block requires the same number of time steps $T$ as the full sequence diffusion, despite that when the block size is small, it may require fewer denoising steps, thus making ACDiT potentially more efficient. Let $\theta$ denote the number of parameters in one layer of the transformer block, $h$ represents the hidden dimension and $n$ as the number of attention heads. The FLOPS is formed by a $O(L)$ component $2L\theta$, from the product between input and weight matrices, and a $O(L^{2})$ component from the pairwise computation between Q-K pairs in the attention dot product, which is $4(h+n)$ for each Q-K pair in the $L^{2}$ Q-K pairs. Consequently, given that $n\ll h$, the total FLOPS can be expressed as: $F=2L\theta+4(h+n)L^{2}\approx 2L\theta+4hL^{2}$.
When employing ACDiT with KV-Cache, the FLOPS consumed by FFN and attention parameters remain unchanged. However, the attention mechanism transforms into causal attention with block size $B$. In this scenario, the number of Q-K pairs to be calculated is given by:
$\sum_{i=1}^{L/B}iB^{2}=\frac{1}{2}(\frac{L}{B}+\frac{L^{2}}{B^{2}})B^{2}.$ (5)
Thus, the FLOPS saved can be expressed as $4h{L^{2}}(\frac{1}{2}(1-\frac{B}{L}))$. Compared to the total FLOPS, the saved percentage is
$\frac{4h{L^{2}}(\frac{1}{2}-\frac{B}{2L})}{4hL^{2}+2L\theta}=\frac{1-B/L}{2+m/k},$ (6)
where $m=\frac{\theta}{h^{2}}$ and $k=\frac{L}{h}$. This equation demonstrates that when the sequence length significantly exceeds the hidden size, transforming a full sequence diffusion into ACDiT can save up to 50% of FLOPS.
However, in practice, it is not always beneficial for small $B$. Setting $B$ to an excessively small value may not fully leverage the iterative modification inherent in the diffusion process, potentially compromising generation quality. Furthermore, given the parallel computing nature of computational kernels, a very small $B$ may not yield speed improvements, a phenomenon analogous to the rationale behind speculative decoding [39]. Conversely, setting $B$ to a very large value diminishes efficiency both in terms of attention calculation, as shown in Equation 6. It also fails to capitalize on the strengths of auto-regressive generation. Indeed, when $B$ is set equal to $L$, ACDiT reverts to the original DiT model. We analyze the influence of block size on performance and efficiency in Sec. 5.3.
Figure 4: The model architecture of ACDiT, mainly inherited from standard DiT architecture. Both the clean latent and noise latent are input while only noise tokens are scale shift with conditioning information. The SCAM attention mask pattern will be applied in the attention block.
Model #Layers Hidden Size MLP #Heads Params
@@ 4.3 Efficiency Analysis and Block Size Choice
Table 1: Configuration details of ACDiT.
[…]
ADM [14] (NeuIPS’19) Diff. - - 554M 10.94 101.0 0.69 0.63
LDM-4-G [58] (CVPR’22) Diff. Cont. - 400M 3.60 247.7 - -
DiT-L/2 [52] (ICCV’23) Diff. Cont. - 458M 5.02 167.2 0.75 0.57
DiT-XL/2 [52] (ICCV’23) Diff. Cont. - 675M 2.27 278.2 0.83 0.57
MaskGIT [76] (CVPR’22) Mask. Disc. - 227M 6.18 316.2 0.83 0.58
MAGE [40] (CVPR’23) Mask. Disc. - 230M 6.93 195.8 - -
VQGAN [21] (CVPR’21) AR Disc. 1.4B 15.78 78.3 - -
RQTran [38] (CVPR’22) AR Disc. 3.8B 7.55 134.0 - -
VAR-d16 [69] (NeurIPS’24) VAR Disc. 310M 3.30 274.4 0.84 0.51
VAR-d20 [69] (NeurIPS’24) VAR Disc. 600M 2.57 302.6 0.83 0.56
LlamaGen-L [67] (arxiv’24) AR Disc. 343M 3.07 256.1 0.83 0.52
LlamaGen-XL [67] (arxiv’24) AR Disc. 775M 2.62 244.1 0.80 0.57
LlamaGen-XXL [67] (arxiv’24) AR Disc. 1.4B 2.34 253.9 0.80 0.59
ImageFolder [42] (ICLR’25) AR Disc. 362M 2.60 295.0 0.75 0.63
LlamaGen+NPP [50] (arxiv’24) AR Cont. 1.4B 2.55 282.0 0.84 0.56
MAR-L [69] (NeurIPS’24) AR Cont. 479M 4.07 232.4 - -
MAR-L [69] (NeurIPS’24) MAR Cont. - 479M 1.78 296.0 0.81 0.60
MonoFormer [86] (arxiv’24) AR + Diff Cont. - 1.1B 2.57 272.6 0.84 0.56
Dart [26] (ICLR’25) AR + Diff Cont. 820M 3.82 263.8 - -
ACDiT-L AR+Diff Cont. 460M 2.53 262.9 0.82 0.55
ACDiT-XL AR+Diff Cont. 677M 2.45 267.4 0.82 0.57
ACDiT-H AR+Diff Cont. 954M 2.37 273.3 0.82 0.57
Table 2: Image generation results on ImageNet 256x256.
4.4 Model Architecture
As shown in Figure 4, ACDiT mainly inherits the main architecture of DiT. We replace the original bidirectional attention with the proposed SCAM attention pattern to process the clean and noise latent. Since we want to keep the architecture as simple and unified as possible, we use linear layers instead of convolution in the input layer and final layer. Besides, we replace the absolute position embedding and Layer Normalization with RoPE [66] and RMSNorm (Root Mean Square Layer Normalization) [70], respectively. We find that QK-norm is important to stabilizing the video generation training, thus we use QK-norms in all experiments. The additional conditional information timesteps and labels are injected into the model with adaLN-Zero only on the noise part. For both image and video generation, we follow DiT and leverage the pre-trained image VAE [35] from Stable Diffusion [58], whose downsample factor is 8. For image generation under B block size, we group square latent representation patches with $\sqrt{B}\times\sqrt{B}$ shape as a block.
5 Experiments
5.1 Experimental Setup
Dataset. For image generation tasks, we consider the ImageNet [60] dataset with 256$\times$256 resolution, which consists of around 1.28M images from 1K classes. For video generation, we consider the UCF-101 [65] dataset with 16 frames, where each frame is an image with 256$\times$256 resolution. UCF-101 contains 13320 videos from 101 classes.
Implementation details. In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. Therefore, for a $256\times 256\times 3$ image in $32\times 32\times 4$ latent shape, the total sequence length and autoregressive length are 1024 and 4, respectively. We explore 4 different model sizes, as shown in Table 1. ACDiT-B is used for design verification and analyze. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer [44] and WSD (Warmup Steady Decay) learning rate scheduling [33] with the peak learning rate 3e-4 and no weight decay. The learning rate begins to decay in the last 15% training iteration. Following the common training recipe of generative models, we keep an exponential moving average (EMA) of the ACDiT weights during training using a decay rate of 0.9999. We sample images with DPM-Solver [45] for 25 steps within each block and use classifier-free guidance [30] with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. For a $16\times 256\times 256\times 3$ video in $16\times 32\times 32\times 4$ latent shape, the sequence length of each frame is 256 and the total sequence length is 4096, with 4 frames grouped into one block. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Other training configs are the same as image training. All models are implemented with PyTorch [51]. Specifically, we use FlexAttention [^3] https://pytorch.org/blog/flexattention to implement the SCAM for both customization and efficiency. We use 256 and 96 NVIDIA H100 GPUs to train the image and video generation model, respectively.
(a) Scaling performance of ACDiT.
(b) Ablation study for ROPE-ND.
(c) FID curve of last 30% training.
Figure 7: (a): ACDiT shows scaling performance similar to DiT. (b): ROPE-ND has consistent improvement to the generation quality. (c): FID score sharply with the learning rate beginning to decay when using the WSD scheduler.
[…]
LVDM [28] (arxiv’22) Diff. 437M 372
Latte [47] (arxiv’24) Diff. 674M 478
Matten [23] (arxiv’24) Diff. 853M 211
VideoFusion [46] (CVPR’23) Diff. 510M 173
MMVG [22] (CVPR’23) Mask. 230M 328
MAGVIT [84] (CVPR’23) Mask. 306M 76
MAGVITv2 [85] (ICLR’24) Mask. 307M 58
TATS [24] (ECCV’22) AR 331M 332
CogVideo [32] (ICLR’23) AR 9.4B 626
MAGVIT-AR [84] (CVPR’23) AR 306M 265
MAGVITv2-AR [85] (ICLR’24) AR 307M 109
OmniTokenizer [74] (NeurIPS’24) AR 650M 191
ACDiT-XL AR+Diff. 677M 111
ACDiT-H AR+Diff. 954M 104
ACDiT-H-LT AR+Diff. 954M 90
Table 3: Video generation results on the UCF-101 dataset. ACDiT-XL-LT means ACDiT-XL trained for longer epoch.
Model Type Top-1 Acc
ViT-H[18] Supervised 83.1
MAGE[40] Masked. 84.3
MAE[27] Masked. 85.9
iGPT [13] Generative 72.6
DiT-XL [52] Generative 82.8
ACDiT-XL Generative 84.0
Table 4: Supervised fine-tuned Top-1 accuracy on Imagenet. We finetune ACDiT-XL and DiT-XL for 100 epoch.
Figure 8: Sample images from ACDiT-H on ImageNet 256$\times$256.
Figure 9: Sample videos from ACDiT-XL trained on UCF-101.
5.2 Main Results
Image Generation. We report the FID-50K [29], Inception Score [61], Precision and Recall [37] of ACDiT and baselines in Table 2. Compared with previous autoregressive models and masked generative models utilizing discrete tokens, such as VQGAN, VAR, LlamaGen, and MaskGIT, ACDiT consistently achieves superior performance with lower FID scores at comparable model scales. Notably, ACDiT-XL achieves 2.45 FID scores, outperforming both LlamaGen-XXL and VAR-d20 with similar parameters. Additionally, when compared to the MAR-L variant that does not recompute attention, ACDiT-L significantly improves performance across all metrics. Although MAR-L has lower FID than ACDiT, recomputing attention makes it hard to generalize longer sequence generation. Compared with other autoregressive diffusion methods, such as Monoformer [86] and Dart [25], ACDiT has a significantly superior performance. When compared with leading diffusion-based methods, ACDiT also demonstrates competitive performance. For instance, despite not employing full-sequence attention, ACDiT models achieve results close to DiT-XL. In general, these results highlight the distinct advantages of ACDiT over other baselines with the continuous latent representation and KV-Cache. Qualitative results are presented in Appendix.
Video Generation. Different from image generation, video inherently includes a temporal dimension, making it more well-suited to autoregressive modeling. The FVD metric on UCF-101 for class-conditional video generation is reported in Table 3. With hybrid AR+Diff architecture, ACDiT-H achieves much lower FVD than other diffusion-based and autoregressive methods, even outperforming MAGVIT-AR and MAGVITv2-AR, which utilizes a closed-source, specially designed video tokenizer. In contrast, ACDiT simplifies the process by directly using an open-sourced image VAE. Although MAGVIT with masked generative methods has a lower FVD than ACDiT, they rely on ”in-place” operation to generate a video similar to the diffusion model. This constraint limits their ability to generalize to generate longer video generation and build world models. Compared to image generation, ACDiT demonstrates a greater potential in the more complex domain of video generation, where longer visual sequences and temporal information are critical. Qualitative results of ACDiT-XL are presented in Fig. 9.
Image Representation. We assess the capability of ACDiT in image representation, which is essential for building a unified visual understanding and generation model. We finetune ACDiT-XL and DiT-XL on ImageNet using classification loss following the training setting in MAE [27] and report the Top-1 accuracy in Table 4. The accuracy of ACDiT surpasses that of ViT-H, iGPT and DiT-XL. This superior performance over DiT-XL highlights the benefit of incorporating clean latent inputs, which accelerates the model’s ability to learn better representations compared to using only noised latent inputs. Furthermore, ACDiT is on par with MAGE [40] in terms of Top-1 accuracy, while enjoying better generation capabilities.

### v3

@@ 4.2 Framework
To satisfy the desiderata discussed above, we propose a versatile framework for autoregressive diffusion called ACDiT. For generality, ACDiT runs block-wise autoregression instead of token-wise autoregression. We identify two kinds of blocks $c_{i}$ and $n_{i}$. For each autoregressive unit $x_{i}$, $c_{i}$ corresponds to the clean version $x_{i}$, $n_{i}$ corresponds to the corrupted version $x_{i}^{(t)}$. ACDiT learns the conditional noise prediction network $\epsilon_{\theta}(n_{i}^{(t)};t,c_{<i})$ with Eq.( 3).
To effectively learn this objective within one transformer network, we can conceptualize all $n_{i}$ and all $c_{i}$ as separate positions, effectively transforming the dependency structure into an attention pattern between different positions. We designate this attention pattern as the Skip-Causal Attention Mask, shown in Figure 2(a). The figure elucidates that $n_{i}$ attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself, while $c_{i}$ also attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself. In training, for simplicity, we can group attention mechanisms as illustrated in the right matrix of Figure 2(a). Suppose the number of blocks is $N$, then the unmasked positions form two triangular matrices with length N1N-1, complemented by a diagonal matrix with length $N$.
During the inference phase, each autoregressive step executes a conditional diffusion process for $n_{i}$ based on $\{c_{0},c_{1},...,c_{i-1}\}$’s KV-Cache. Upon finishing denoising, it is appended to the clean sequence as $c_{i}$ followed by the maximal noisy version of the next block $n_{i+1}$. The key-value tensor will be computed for these two blocks, and the key-value tensor of the clean block $c_{i}$ will be kept in KV-Cache. All noisy version of $n_{i}$ is disregarded. The process is visualized in Figure 2(b). A three-dimensional view is presented in Figure 2(c). By interpolating between full-sequence diffusion and autoregressive paradigms, ACDiT enjoys flexibility and expressivity, enabling it to generate video of any length utilizing the latest long-context techniques developed for language models.
(a) Comparison of model design.
(b) The model architecture of ACDiT.
Figure 3: (a): Comparison between ACDiT with Transfusion (Zhou et al., 2024a), Diffusion Forcing (Chen et al., 2024a) and MAR (Li et al., 2024a).Transfusion does not perform autoregressive modeling within the visual modality. Diffusion Forcing introduces different noise levels across the autoregressive inputs. MAR applies diffusion only on top of the backbone model. ACDiT utilizes the full parameters for both the autoregressive and diffusion process with clean input as context. (b): The model architecture of ACDiT. Both the clean and noisy latent are input, while only the noisy latent interacts with the conditioning information. The Skip-Causal Attention Mask pattern, shown in Figure 2(a), will be applied in the attention block.
@@ 4.3 Positional Encoding
ACDiT is designed to be versatile, capable of handling one-, two-, three-, or even higher-dimensional data, including but not limited to text (1D), images (2D), and video (3D). For any given dimension of data, the position of that data is a critical attribute that must be made known to the model. This positional awareness enables the model to contextualize its current focus relative to historical data. In the domain of textual data, the Rotary Position Embedding (RoPE) (Su et al., 2024) has gained widespread adoption as an effective relative positional encoding method. To address the challenges posed by multi-dimensional positional indices, we introduce RoPE-ND, a natural extension of RoPE.
For a token of a $D$ dimensional data, its positional index is $m_{1},m_{2},...,m_{D}]$. Given query and key vectors in the Transformer’s attention module, we partition the hidden dimension into $D$ segments. It is imperative that each segment’s hidden dimension be an even number. For each segment $j$, we apply a RoPE with a specific base $b_{j}$, as defined in the following equation:
$\mathbf{R}=\begin{bmatrix}\mathbf{R}^{d_{1}}_{\Theta_{1},m_{1}}&0&\cdots&0\\ 0&\mathbf{R}^{d_{2}}_{\Theta_{2},m_{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\mathbf{R}^{d_{D}}_{\Theta_{D},m_{D}}\end{bmatrix}$, (4)
where each $\mathbf{R}^{d_{j}}_{\Theta_{j},m_{j}}$ represents a $d_{j}$-dimensional rotary matrix [^2] See a detailed explanation in the Equation 15 in  (Su et al., 2024). with rotation angles $\Theta_{j}=\{b_{j}^{-2(i-1)/d_{j}},i\in[1,2,\cdots,\frac{d_{j}}{2}]\}$. The base $b_{j}$ is empirically determined as $100\lceil\frac{8L_{j}}{100\pi}\rceil$, where $L_{j}$ denotes the maximum position index in dimension $j$. This formulation ensures that the highest wavelength of RoPE is approximately eight times the maximum position, thereby mitigating rapid decay in long-term dependencies. It is worth noting that ACDiT inherently supports length extrapolation (Su et al., 2024), although a comprehensive exploration of this falls beyond the present work.
4.4 Efficiency Analysis and Block Size Choice
We provide a brief analysis of the computational efficiency of ACDiT in terms of floating-point operations (FLOPS). The full derivation and more analysis are discussed in the Appendix A. We first assume that denoising each block requires the same number of time steps $T$ as the full sequence diffusion, despite that when the block size is small, it may require fewer denoising steps, thus making ACDiT potentially more efficient. Let $L$ denote the sequence length, $h$ the hidden dimension, $\theta$ the number of parameters in a transformer layer, and $B$ the block size used in ACDiT. The FLOPS for standard full-sequence diffusion is $F_{\text{full}}=2L\theta+4hL^{2}$. With KV-Cache, ACDiT reduces the attention cost to $2hL^{2}+2hLB$. The relative FLOPS savings is: $\frac{F_{\text{saved}}}{F_{\text{full}}}=\frac{1-\tfrac{B}{L}}{2+\tfrac{\theta}{hL}}.$ This indicates that ACDiT can reduce per-layer FLOPS by up to 50% when $L\gg B$.
Table 1: Image generation results on ImageNet 256x256. We report FID score, Inception Score (IS), Precision (Pre.), and Recall (Rec.).
@@ 4.4 Efficiency Analysis and Block Size Choice
ADM (Dhariwal and Nichol, 2021) Diff. - - 554M 10.94 101.0 0.69 0.63
LDM-4-G (Rombach et al., 2022) Diff. Cont. - 400M 3.60 247.7 - -
DiT-L/2 (Peebles and Xie, 2022) Diff. Cont. - 458M 5.02 167.2 0.75 0.57
DiT-XL/2 (Peebles and Xie, 2022) Diff. Cont. - 675M 2.27 278.2 0.83 0.57
MaskGIT (Weber et al., 2024) Mask. Disc. - 227M 6.18 316.2 0.83 0.58
MAGE (Li et al., 2023) Mask. Disc. - 230M 6.93 195.8 - -
VQGAN (Esser et al., 2021) AR Disc. 1.4B 15.78 78.3 - -
RQTran (Lee et al., 2022) AR Disc. 3.8B 7.55 134.0 - -
VAR-d16 (Tian et al., 2024) VAR Disc. 310M 3.30 274.4 0.84 0.51
VAR-d20 (Tian et al., 2024) VAR Disc. 600M 2.57 302.6 0.83 0.56
LlamaGen-L (Sun et al., 2024) AR Disc. 343M 3.07 256.1 0.83 0.52
LlamaGen-XL (Sun et al., 2024) AR Disc. 775M 2.62 244.1 0.80 0.57
LlamaGen-XXL (Sun et al., 2024) AR Disc. 1.4B 2.34 253.9 0.80 0.59
ImageFolder (Li et al., 2024b) AR Disc. 362M 2.60 295.0 0.75 0.63
LlamaGen+NPP (Pang et al., 2024) AR Cont. 1.4B 2.55 282.0 0.84 0.56
MAR-L-Causal (Li et al., 2024a) AR Cont. 479M 4.07 232.4 - -
MAR-L (Li et al., 2024a) Mask. Cont. - 479M 1.78 296.0 0.81 0.60
MonoFormer (Zhao et al., 2024a) AR+Diff Cont. - 1.1B 2.57 272.6 0.84 0.56
Dart (Gu et al., 2025) AR+Diff Cont. 820M 3.82 263.8 - -
ACDiT-L AR+Diff Cont. 460M 2.53 262.9 0.82 0.55
ACDiT-XL AR+Diff Cont. 677M 2.45 267.4 0.82 0.57
ACDiT-H AR+Diff Cont. 954M 2.37 273.3 0.82 0.57
4.5 Model Architecture and Implementation
As shown in Fig. 3(b), ACDiT mainly inherits the main architecture of DiT. We replace the original bidirectional attention with the proposed SCAM attention pattern to process the clean and noisy latent. In addition, we replace the absolute position embedding and Layer Normalization with RoPE (Su et al., 2024) and RMSNorm (Touvron et al., 2023a), respectively. We use QK-norms for stable training. The additional condition, including timesteps and labels, is injected into the model with AdaLN-Zero only on the noise part.
We explore 4 different model sizes, as shown in Table 6. In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer (Loshchilov, 2017) and WSD (Warmup Steady Decay) learning rate scheduling (Hu et al., 2024). We sample images with DPM-Solver (Lu et al., 2022) for 25 steps within each block and use classifier-free guidance (Ho and Salimans, 2022) with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Unless otherwise specified, ablation and analysis experiments are conducted on ACDiT-B. See more details at Appendix B.
5 Experiments
We mainly experiment on the ImageNet (Russakovsky et al., 2015) dataset with $\times256$ resolution and the UCF-101 (Soomro et al., 2012) dataset with 16 frames for image generation and video generation, respectively. In addition, we also validate the generality of ACDiT on text generation.
5.1 Main Results
Image Generation. We report the FID-50K (Heusel et al., 2017), Inception Score (Salimans et al., 2016), Precision and Recall (Kynkäänniemi et al., 2019) of ACDiT and baselines in Table 1. Compared with previous autoregressive models and masked generative models utilizing discrete tokens, such as VQGAN, VAR, LlamaGen, and MaskGIT, ACDiT consistently achieves superior performance with lower FID scores at comparable model scales. Notably, ACDiT-XL achieves 2.45 FID scores, outperforming both LlamaGen-XXL and VAR-d20 with similar parameters. Additionally, when compared to the MAR-L-Causal variant that does not recompute attention, ACDiT-L significantly improves performance across all metrics, specifically improving FID from 4.07 to 2.53. Compared with other autoregressive diffusion methods, such as Monoformer (Zhao et al., 2024a) and Dart (Gu et al., 2024), ACDiT has a significantly superior performance. When compared with leading diffusion-based methods, ACDiT also demonstrates competitive performance. For instance, despite not employing full-sequence attention, ACDiT models achieve results close to DiT-XL. In general, these results highlight the distinct advantages of ACDiT over other baselines with the continuous latent representation and KV-Cache. Qualitative results are presented in the Appendix E.
Table 2: Video generation results on the UCF-101 dataset. ACDiT-XL-LT means ACDiT-XL trained for longer epochs.
[…]
LVDM (He et al., 2022b) Diff. 437M 372
Latte (Ma et al., 2024) Diff. 674M 478
Video-LaVIT (Jin et al., 2024) Diff. 7B 281
MagDiff (Zhao et al., 2024b) Diff. 2B 340
Matten (Gao et al., 2024) Diff. 853M 211
VideoFusion (Luo et al., 2023) Diff. 510M 173
MMVG (Fu et al., 2023) Mask. 230M 328
MAGVIT (Yu et al., 2023a) Mask. 306M 76
MAGVITv2 (Yu et al., 2023b) Mask. 307M 58
TATS (Ge et al., 2022) AR 331M 332
CogVideo (Hong et al., 2022) AR 9.4B 626
MAGVIT-AR (Yu et al., 2023a) AR 306M 265
MAGVITv2-AR (Yu et al., 2023b) AR 307M 109
OmniTokenizer (Wang et al., 2024a) AR 650M 191
ACDiT-XL AR+Diff. 677M 111
ACDiT-H AR+Diff. 954M 104
ACDiT-H-LT AR+Diff. 954M 90
Table 3: Classification accuracy on ImageNet.
Model Type Params Top-1 Acc
ViT-H (Dosovitskiy, 2020) Supervised 632M 83.1
MAGE (Li et al., 2023) Masked 328M 84.3
MAE (He et al., 2022a) Masked 632M 85.9
iGPT (Chen et al., 2020) Generative 1.4B 72.6
DiT-XL (Peebles and Xie, 2022) Generative 675M 82.8
ACDiT-XL Generative 677M 84.0
Table 4: Test perplexities on OpenWebText.
PPL ($\downarrow$)
Autoregressive
AR (Sahoo et al., 2024) 17.54
Diffusion
SEDD (Lou et al., 2023) $\leq$ 24.10
MDLM (Sahoo et al., 2024) $\leq$ 22.98
Block-wise diffusion
BD3-LMs (Arriola et al., 2025) $L^{\prime}=16$ $\leq$ 22.27
                  $L^{\prime}=8$ $\leq$ 21.68
ACDiT            $L^{\prime}=256$ $\leq$ 22.98
                  $L^{\prime}=128$ $\leq$ 22.82
                  $L^{\prime}=8$ $\leq$ 21.59
Video Generation. Different from image generation, video inherently includes a temporal dimension, making it more well-suited to autoregressive modeling. The FVD metric on UCF-101 for class-conditional video generation is reported in Table 2. With hybrid AR+Diff architecture, ACDiT-H achieves much lower FVD than other diffusion-based and autoregressive methods, even outperforming MAGVIT-AR and MAGVITv2-AR, which utilize a closed-source, specially designed video tokenizer. In contrast, ACDiT simplifies the process by directly using an open-sourced image VAE. Although MAGVIT with masked generative methods has a lower FVD than ACDiT, they rely on in-place operation to generate a video similar to the diffusion model and require repeatedly processing global context during inference. This constraint limits their ability and efficiency to generalize to long video generation and build world models. We further analyze the efficiency implications of autoregressive versus masked generative formulations in the following section, with quantitative inference-time comparisons reported in Table 6. Compared to image generation, ACDiT demonstrates a greater potential in the more complex domain of video generation, where longer visual sequences and temporal information are critical. Qualitative results are presented in the Appendix E.
Text Generation. ACDiT is a general, modality-agnostic framework. At the architectural level, ACDiT adopts a unified block-wise autoregressive formulation and a shared attention pattern that conditions each block on clean past context while allowing non-causal modeling within the current block. This design is independent of the data modality and remains unchanged when applied to text generation. For text generation, which operates on discrete tokens, the differences from visual generation arise only in the noise process used during training and inference. Specifically, we follow discrete diffusion language modeling objectives and the same experimental setup as previous work (Sahoo et al., 2024; Shi et al., 2024). Importantly, these choices affect only the modality-specific corruption and denoising operators, while the model architecture, attention mask, and block-wise autoregressive structure remain identical. Specifically, we use a 12-layer Transformer with a hidden dimension of 768, 12 attention heads, and a 128-dimensional timestep embedding as our model backbone. We conduct all training and validation on the OpenWebText dataset (Gokaslan and Cohen, 2019), an open-source replica of the WebText dataset. To evaluate the model’s ability in probabilistic modeling and compression, we report Perplexity (PPL) on the validation set. PPL is defined as:
$\mathrm{PPL}=\exp\left(\frac{\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}}}\left[-\log p_{\theta}(\bm{x}_{0})\right]}{D}\right),$
where $D$ is the data dimension. For models where the exact likelihood $p_{\theta}(\bm{x}_{0})$ is intractable, we report the PPL calculated using the variational lower bound on the log-likelihood. Despite conceptual similarity to blockwise discrete approaches (Arriola et al., 2025), our effective positional embedding implementation (e.g., RoPE-ND; Sec. 4.3) enables ACDiT to achieve comparable performance across various block sizes as shown in Tab. 4 without any customized variance reduction or noise schedules as proposed in (Arriola et al., 2025). These advantages highlight the potential of ACDiT as a general and powerful generative modeling framework.
Image Representation. We assess the capability of ACDiT in image representation, which is essential for building a unified visual understanding and generation model. We finetune ACDiT-XL and DiT-XL on ImageNet using classification loss for 100 epochs and report the Top-1 accuracy in Table 4. ACDiT outperforms DiT-XL, highlighting the advantage of using clean latents, which help the model to learn better representations compared to using only noisy latent inputs. Additionally, ACDiT matches the Top-1 accuracy of MAGE (Li et al., 2023), while offering superior generation performance.
(a) Scaling performance of ACDiT.
(b) Ablation for ROPE-ND and SCAM.
(c) FID curve of last 30% training.
Figure 6: (a): ACDiT shows scaling performance similar to DiT. (b): Ablation for ROPE-ND and using clean contexts. (c): FID score sharply drops with the learning rate decaying when using the WSD scheduler.
Table 5: Configuration details of ACDiT.
Model #Layers Hidden MLP #Heads Params
@@ 5.1 Main Results
Table 6: Inference cost comparison with MAR.
Model Params Image Video
MAR-B 322M 2.74s 49.27s
MAR-L 479M 3.58s 63.82s
ACDiT-L 460M 2.29s 4.08s
ACDiT-XL 677M 3.22s 5.16s